• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 09:20:23 on March 30, 2008

    Entry number 229030

    keyword=Update on beam delivery

    The machine had a vaccum problem yesterday and it was decided to redo the optics which was very unstable. Today will be dedicated to retune the machine so most likely beam will be ready to be delivered by monday morning. We will lock-up when MCC will be ready to deliver beam in Hall A. So most likely we will stay in restricted access until monday unless I get an update from the machine that they are ready to send beam in the Hall earlier.